# On $L^{p}$ extremals for Fourier extension estimate to fractional surface

### v1

@@ 4. The $L^{p}$ convergence
Throughout this section, we assume the extension operator $\mathcal{E}_{\alpha}$ extends as a bounded operator from $L^{p_{0}}(\mathbb{R}^{d})$ to $L^{q_{0}}(\mathbb{R}^{d+1})$ for some $1<p_{0}<q_{0}=\frac{d+\alpha}{d}p_{0}^{\prime}$. We investigate the case $1<p<p_{0}$ with $q=\frac{d+\alpha}{d}p^{\prime}$. In this section, we show the desired $L^{p}$-precompactness result Theorem 1.1 by following the outline given by Stovall [47].
@@ 4.1. Frequency localization
We first demonstrate that extremal sequences possess good frequency localization, which relies one some refinements of the original Strichartz estimate (1).
@@
For the function $f\in L^{p}(\mathbb{R}^{d})$, there exists a parameter $\theta_{0}\in(0,1)$ such that
$\|\mathcal{E}_{\alpha}f\|_{L^{q}(\mathbb{R}^{d+1})}\lesssim\sup_{N\in\mathbb{Z}}\|\mathcal{E}_{\alpha}f_{\mathcal{A}_{N}}\|_{L^{q}(\mathbb{R}^{d+1})}^{\theta_{0}}\|f\|_{L^{p}(\mathbb{R}^{d})}^{1-\theta_{0}}.$
@@
We adapt some ideas from [38, Proof of Lemma 3.2]. By Lemma 2.9, we only need to show the following bilinear estimate
@@
Without loss of generality, we may assume $N_{1}>N_{2}$. Choose $(r_{1},q_{1},s_{1})$ and $(r_{2},q_{2},s_{2})$ satisfying
$\dfrac{d}{r_{1}}+\dfrac{2}{q_{1}}\leq\dfrac{d}{2},\quad\dfrac{d}{r_{2}}+\dfrac{2}{q_{2}}\leq\dfrac{d}{2},\quad\frac{d}{d+\alpha}=\frac{1}{q_{1}}+\frac{1}{q_{2}}=\frac{1}{r_{1}}+\frac{1}{r_{2}}.$
@@
$\lesssim$ $2^{s_{2}(N_{1}-N_{2})}\|f_{\mathcal{A}_{N_{1}}}\|_{L^{2}(\mathbb{R}^{d})}\|f_{\mathcal{A}_{N_{2}}}\|_{L^{2}(\mathbb{R}^{d})}.$
Further choose suitable $\widetilde{p}\leq p_{0}$ so that $p$ lies between $2$ and $\widetilde{p}$. Then, applying Hölder’s inequality and the assumption on $p_{0}$, we obtain
$\|\mathcal{E}_{\alpha}f_{\mathcal{A}_{N_{1}}}\mathcal{E}_{\alpha}f_{\mathcal{A}_{N_{2}}}\|_{L^{\frac{\widetilde{q}}{2}}(\mathbb{R}^{d+1})}\lesssim\|f_{\mathcal{A}_{N_{1}}}\|_{L^{\widetilde{p}}(\mathbb{R}^{d})}\|f_{\mathcal{A}_{N_{2}}}\|_{L^{\widetilde{p}}(\mathbb{R}^{d})},\quad\widetilde{q}:=\frac{(d+\alpha)\widetilde{p}^{\prime}}{d}.$
By interpolating the aforementioned two estimates, we obtain the desired bilinear estimate and complete the proof. ∎
@@
For each $f\in L^{p}(\mathbb{R}^{d})$, there exist $\theta\in(0,1)$ and $c_{0}>0$ such that
(25) $\|\mathcal{E}_{\alpha}f\|_{L^{q}(\mathbb{R}^{d+1})}\lesssim\sup\limits_{k\in\mathbb{Z}}\sup\limits_{\tau\in\mathcal{D}_{k}}\sup\limits_{l\geq 0}2^{-c_{0}l}\|f_{\tau,l}\|_{L^{p}(\mathbb{R}^{d})}^{\theta}\|f\|_{L^{p}(\mathbb{R}^{d})}^{1-\theta},$
@@
For $q_{1}>\frac{2(d+3)}{d+1}$, Tao’s bilinear restriction estimate (see Lemma 2.2) implies that
$\|\mathcal{E}_{\alpha}f_{\tau}\mathcal{E}_{\alpha}f_{\tau^{\prime}}\|_{L^{\frac{q_{1}}{2}}(\mathbb{R}^{d+1})}\lesssim|\tau|^{1-2\frac{d+2}{dq_{1}}}\|f_{\tau}\|_{L^{2}(\mathbb{R}^{d})}\|f_{\tau^{\prime}}\|_{L^{2}(\mathbb{R}^{d})}\lesssim|\tau|^{1-2\frac{d+\alpha}{dq_{1}}}\|f_{\tau}\|_{L^{2}(\mathbb{R}^{d})}\|f_{\tau^{\prime}}\|_{L^{2}(\mathbb{R}^{d})}$
holds for all $f\in L^{2}(\mathbb{R}^{d})$ with $\mathrm{supp}f\subset\mathcal{A}_{0}$ and for all $\tau\sim\tau^{\prime}\subset\mathcal{A}_{0}$. Here the second inequality is followed by $|\tau|\leq 1$ and $\alpha>2$. Then for $1<p_{2}<p_{0}$ and $q_{2}=\frac{d+\alpha}{d}p_{2}^{\prime}$, by applying Hölder’s inequality and the assumption on $p_{0}$, we can then use rescaling to conclude that
$\|\mathcal{E}_{\alpha}f_{\tau}\mathcal{E}_{\alpha}f_{\tau^{\prime}}\|_{L^{\frac{q_{2}}{2}}(\mathbb{R}^{d+1})}\lesssim\|f_{\tau}\|_{L^{p_{2}}(\mathbb{R}^{d})}\|f_{\tau^{\prime}}\|_{L^{p_{2}}(\mathbb{R}^{d})}.$
Interpolating these two bilinear estimates with suitable $q_{1}$ and $q_{2}$, there exists $s<p$ such that
@@
$\left\|\mathcal{E}_{\alpha}f_{\mathcal{A}_{0}^{j}}\right\|_{L^{q}(\mathbb{R}^{d+1})}^{2}\leq\left\|\sum_{\tau\sim\tau^{\prime}\subset\mathcal{A}_{0}^{j}}\mathcal{E}_{\alpha}f_{\tau}\mathcal{E}_{\alpha}f_{\tau^{\prime}}\right\|_{L^{\frac{q}{2}}(\mathbb{R}^{d+1})}\lesssim\left(\sum_{\tau\sim\tau^{\prime}\subset\mathcal{A}_{0}^{j_{0}}}\left\|\mathcal{E}f_{\tau}\mathcal{E}f_{\tau^{\prime}}\right\|_{L^{\frac{q}{2}}(\mathbb{R}^{d+1})}^{t}\right)^{\frac{1}{t}},$
where $t:=(q/2)_{*}>p/2>s/2$ and recall $(q/2)_{*}:=\min\{q/2,(q/2)^{\prime}\}$. By (26), we further have
@@
$\left\|\mathcal{E}_{\alpha}f_{\mathcal{A}_{0}^{j}}\right\|_{L^{q}(\mathbb{R}^{d+1})}^{2}\lesssim\left(\sum_{k}\sum_{\tau\in\mathcal{D}_{k}}|\tau|^{-2t(\frac{1}{s}-\frac{1}{p})}\|f_{\tau}\|_{L^{s}(\mathbb{R}^{d})}^{2t}\right)^{\frac{1}{t}}.$
@@
$\lesssim$ $\left(\sum_{k}\sum_{\tau\in\mathcal{D}_{k}}|\tau|^{-2t(\frac{1}{s}-\frac{1}{p})}\|f_{\tau}\|_{L^{s}(\mathbb{R}^{d})}^{2t}\right)^{\frac{1}{2t}}$
$\lesssim$ $\sup_{k\in\mathbb{Z}}\sup_{\tau\in\mathcal{D}_{k}}\sup_{l\geq 0}2^{-c_{1}l}\|f_{\tau,l}\|_{L^{p}(\mathbb{R}^{d})}^{\theta_{1}}\|f\|_{L^{p}(\mathbb{R}^{d})}^{1-\theta_{1}}$
for some $c_{1}>0$ and $10<\theta_{1}<$ 1. Then by rescaling, we obtain
$\left\|\mathcal{E}_{\alpha}f_{\mathcal{A}_{N}}\right\|_{L^{q}(\mathbb{R}^{d+1})}\lesssim\sup_{k\in\mathbb{Z}}\sup_{\tau\in\mathcal{D}_{k}}\sup_{l\geq 0}2^{-c_{1}l}\|f_{\tau,l}\|_{L^{p}(\mathbb{R}^{d})}^{\theta_{1}}\|f\|_{L^{p}(\mathbb{R}^{d})}^{1-\theta_{1}}$
holds for all $f\in L^{p}(\mathbb{R}^{d})$ with $\mathrm{supp}f\subset\mathcal{A}_{N}$. Based on this estimate and Lemma 4.1, we get the desired estimate (25) and finish the proof. ∎
@@
We are going to show that the desired decreasing sequence can be taken as $\rho_{j}:=C_{*}(j/2)^{-\theta/p}$, where $\theta$ is taken from Lemma 4.2 and $C_{*}$ is the optimal constant of the inequality (25). Then for fixed $j$, we investigate the following item
$\sup\limits_{\tau\in\mathcal{D}}\|f\chi_{\tau}\chi_{\{\xi:|f(\xi)|<2^{j}|\tau|^{-{1}/{p}}\|f\|_{L^{p}(\mathbb{R}^{d})}\}}\|_{L^{p}(\mathbb{R}^{d})}\neq 0.$
Let the cube sequence $\tau_{n}=c_{n}+h_{n}[0,1]^{d}$ approach this supremum (bigger than half of this supremum). Then one can use dominated convergence theorem to show that the cases $h_{n}\to\infty$ and $h_{n}\to 0$ cannot occur, and moreover, the case $c_{n}\to\infty$ cannot occur. These imply the fact that the aforementioned supremum can be reached.
@@
One can observe that the functions $f^{j}$ in Lemma 4.3, even with small support, may have large values, and conversely, $f^{j}$ with small values may have wide support. Now, we aim to control both the support and the value simultaneously.
@@
If this proposition were to fail, for any dilation parameters $\left\{(h_{n})\right\}\subset\mathbb{R}_{+}$, there would exist $\varepsilon>0$ and $\lim_{l\to\infty}m_{l}=\infty$ such that
@@
Recalling that there exists $J_{*}$ such that $\rho_{J_{*}}<M_{\alpha,p}$, then up to subsequences, for some $j\in\{1,2,\ldots,J_{*}\}$ there holds
(31) $\|\mathcal{E}_{\alpha}f_{n}^{j}\|_{L^{q}(\mathbb{R}^{d+1})}\geq\frac{M_{\alpha,p}-\rho_{J_{*}}}{2J_{*}}>0,\quad\forall n\in\mathbb{N}.$
Without loss of generality, we may assume $j=1$. Let
$\tau_{n}^{1}:=\xi_{n}^{1}+2^{-k_{n}^{1}}[0,1]^{d},\quad g_{n}^{1}(\xi):=2^{\frac{k_{n}^{1}d}{p}}f_{n}^{1}\left(2^{-k_{n}^{1}}\xi\right),\quad\Rightarrow\quad\left|g_{n}^{1}\left(\xi-2^{k_{n}^{1}}\xi_{n}^{1}\right)\right|\leqslant 2\chi_{[0,1]^{d}}(\xi)$
If there holds $\lim\limits_{n\to\infty}2^{k_{n}^{1}}|\xi_{n}^{1}|=\infty$, similar to the proof of the stationary phase consequence (6) , then the fact $\alpha>2$ and dominated convergence theorem imply that the following item goes to zero as $n$ goes to infinity
@@
$=\left(2^{k_{n}^{1}}|\xi_{n}^{1}|\right)^{2-\alpha}\int_{\mathbb{R}^{d+1}}\left|\int_{\mathbb{R}^{d}}e^{ix\xi+it\frac{|\xi-2^{k_{n}^{1}}\xi_{n}^{1}|^{\alpha}-\left(2^{k_{n}^{1}}|\xi_{n}^{1}|\right)^{\alpha}+\alpha\left(2^{k_{n}^{1}}|\xi_{n}^{1}|\right)^{\alpha-2}2^{k_{n}^{1}}\xi\cdot\xi_{n}^{1}}{\left(2^{k_{n}^{1}}|\xi_{n}^{1}|\right)^{\alpha-2}}}g_{n}^{1}\left(\xi-2^{k_{n}^{1}}\xi_{n}^{1}\right)\mathrm{d}\xi\right|^{q}\mathrm{d}x\mathrm{d}t,$
which contradicts the inequality (31). Thus up to subsequences, $2^{k_{n}^{1}}\xi_{n}^{1}\to\xi^{1}$ as $n\to\infty$.
By applying scaling symmetries, we can assume $\tau_{n}^{1}=\xi^{1}+[0,1]^{d}$ for all $n$ and the remainder cubes can be written as
@@
By hypothesis, for such $\{h_{n}\}$ we get $\varepsilon$ and $m_{l}$ satisfying inequality (29), and by taking $J$ large enough the inequality (30) implies that
@@
Next we are going to apply the generalized Brézis-Lieb lemma [22, Lemma 3.1], so we need to show that $\mathcal{E}_{\alpha}B_{n}^{J}(x,t)\to 0$ almost everywhere in $\mathbb{R}^{d+1}$ as $n\to\infty$. Since $J$ is a fixed finite number, we only need to investigate each item $\mathcal{E}_{\alpha}f_{n}^{j}$ with bad index $j$. For each fixed bad index $j$, we can divide the proof into three cases: if $k_{n}^{j}\to+\infty$ as $n$ goes to infinity then $f_{n}^{j}\to 0$ in $L^{1}(\mathbb{R}^{d})$, which by direct computation implies $\mathcal{E}_{\alpha}f_{n}^{j}\to 0$ almost everywhere; if $k_{n}^{j}\to-\infty$ as $n$ goes to infinity then $f_{n}^{j}\to 0$ in $L^{p_{0}}(\mathbb{R}^{d})$, which by the assumption on $p_{0}$ further implies $\mathcal{E}_{\alpha}f_{n}^{j}\to 0$ in $L^{q_{0}}(\mathbb{R}^{d+1})$ and up to subsequences $\mathcal{E}_{\alpha}f_{n}^{j}\to 0$ almost everywhere; if $|k_{n}^{j}|$ remains bounded and $|\xi_{n}^{j}|\to\infty$ as $n$ goes to infinity, which implies $f_{n}^{j}\rightharpoonup 0$ in $L^{2}(\mathbb{R}^{d})$ and further the Lemma 2.4 implies $\mathcal{E}_{\alpha}f_{n}^{j}(t,x)\to 0$ almost everywhere. In summary, for fixed $J$, we obtain the desired pointwise convergence conclusion $\mathcal{E}_{\alpha}B_{n}^{J}\to 0$ almost everywhere. By the construction, for fixed $J$, there exists $G^{J}\in L^{p}(\mathbb{R}^{d})$ such that $G_{n}^{J}\to G^{J}$ in $L^{p}(\mathbb{R}^{d})$ as $n$ goes to infinity. Hence, by the aforementioned generalized Brézis-Lieb lemma and the fact $q>p$, we conclude

### v2

@@ 4. The $L^{p}$ convergence
In this section, we show the desired $L^{p}$-precompactness result Theorem 1.1. For the case $p=1$, one can see that $M_{\alpha,1}=1$, hence the existence of extremals and the non-compactness of extremal sequences are elementary. Recall that the typical $p=2$ case has been solved in Corollary 3.6. Thus throughout this section, we investigate the non-endpoint case $p\in(1,2)$ with $q=\frac{d+\alpha}{d}p^{\prime}$.
In Subsection 4.1, we establish some $L^{p}$-refined Strichartz estimates and show some frequency localization properties for the extremal sequences; then in Subsection 4.2, we apply the $L^{2}$-theory established in the previous section to show some space-time localization properties for the extremal sequences; finally in Subsection 4.3, we show the desired $L^{p}$-precompactness Theorem 1.1.
@@ 4.1. Frequency localization
We first demonstrate that extremal sequences possess good frequency localization, which relies on some refinements of the original Strichartz estimate (1).
@@
There exists a parameter $\theta_{0}\in(0,1)$ such that
$\|\mathcal{E}_{\alpha}f\|_{L^{q}(\mathbb{R}^{d+1})}\lesssim\sup_{N\in\mathbb{Z}}\|\mathcal{E}_{\alpha}f_{\mathcal{A}_{N}}\|_{L^{q}(\mathbb{R}^{d+1})}^{\theta_{0}}\|f\|_{L^{p}(\mathbb{R}^{d})}^{1-\theta_{0}},\quad\forall f\in L^{p}(\mathbb{R}^{d}).$
@@
We adapt some ideas from [33, Proof of Lemma 3.2], which investigates the cone case. By Lemma 2.9, it suffices to show that there exists $c_{1}>0$ and the following bilinear estimate
@@
Without loss of generality, we may assume $N_{1}>N_{2}$. Then we choose two pairs $(r_{1},q_{1},s_{1})$ and $(r_{2},q_{2},s_{2})$ which satisfy the following relations
$\dfrac{d}{r_{1}}+\dfrac{2}{q_{1}}\leq\dfrac{d}{2},\quad\dfrac{d}{r_{2}}+\dfrac{2}{q_{2}}\leq\dfrac{d}{2},\quad\frac{d}{d+\alpha}=\frac{1}{q_{1}}+\frac{1}{q_{2}}=\frac{1}{r_{1}}+\frac{1}{r_{2}},$
@@
$\lesssim$ $2^{-s_{2}(N_{1}-N_{2})}\|f_{\mathcal{A}_{N_{1}}}\|_{L^{2}(\mathbb{R}^{d})}\|f_{\mathcal{A}_{N_{2}}}\|_{L^{2}(\mathbb{R}^{d})}.$
Recalling that $\mathcal{E}_{\alpha}$ is $L^{1}(\mathbb{R}^{d})$ to $L^{\infty}(\mathbb{R}^{d+1})$ bounded, we obtain
$\|\mathcal{E}_{\alpha}f_{\mathcal{A}_{N_{1}}}\mathcal{E}_{\alpha}f_{\mathcal{A}_{N_{2}}}\|_{L^{\infty}(\mathbb{R}^{d+1})}\lesssim\|f_{\mathcal{A}_{N_{1}}}\|_{L^{1}(\mathbb{R}^{d})}\|f_{\mathcal{A}_{N_{2}}}\|_{L^{1}(\mathbb{R}^{d})}.$
Finally, by interpolating the aforementioned two estimates on $L^{2}$ and $L^{1}$, we obtain the desired bilinear estimate and complete the proof. ∎
@@
There exist $\theta\in(0,1)$ and $c_{0}>0$ such that
(25) $\|\mathcal{E}_{\alpha}f\|_{L^{q}(\mathbb{R}^{d+1})}\lesssim\sup\limits_{k\in\mathbb{Z}}\sup\limits_{\tau\in\mathcal{D}_{k}}\sup\limits_{l\geq 0}2^{-c_{0}l}\|f_{\tau,l}\|_{L^{p}(\mathbb{R}^{d})}^{\theta}\|f\|_{L^{p}(\mathbb{R}^{d})}^{1-\theta},\quad\forall f\in L^{p}(\mathbb{R}^{d}),$
@@
For arbitrary $q_{1}>\frac{2(d+3)}{d+1}$, Tao’s bilinear restriction estimate (see Lemma 2.2) implies that
(26) $\|\mathcal{E}_{\alpha}f_{\tau}\mathcal{E}_{\alpha}f_{\tau^{\prime}}\|_{L^{\frac{q_{1}}{2}}(\mathbb{R}^{d+1})}\lesssim|\tau|^{1-2\frac{d+2}{dq_{1}}}\|f_{\tau}\|_{L^{2}(\mathbb{R}^{d})}\|f_{\tau^{\prime}}\|_{L^{2}(\mathbb{R}^{d})}\lesssim|\tau|^{1-2\frac{d+\alpha}{dq_{1}}}\|f_{\tau}\|_{L^{2}(\mathbb{R}^{d})}\|f_{\tau^{\prime}}\|_{L^{2}(\mathbb{R}^{d})}$
holds for all $f\in L^{2}(\mathbb{R}^{d})$ with $\mathrm{supp}f\subset\mathcal{A}_{0}$ and for all $\tau\sim\tau^{\prime}\subset\mathcal{A}_{0}$. Here the second inequality is followed by $|\tau|\leq 1$ and $\alpha>2$. For arbitrary $p_{2}\in[1,2]$ and $q_{2}=\frac{d+\alpha}{d}p_{2}^{\prime}$, $\mathcal{E}_{\alpha}$ is bounded from $L^{p}(\mathbb{R}^{d})$ to $L^{q}(\mathbb{R}^{d+1})$. Thus by Hölder’s inequality, we can conclude that
(27) $\|\mathcal{E}_{\alpha}f_{\tau}\mathcal{E}_{\alpha}f_{\tau^{\prime}}\|_{L^{\frac{q_{2}}{2}}(\mathbb{R}^{d+1})}\lesssim\|f_{\tau}\|_{L^{p_{2}}(\mathbb{R}^{d})}\|f_{\tau^{\prime}}\|_{L^{p_{2}}(\mathbb{R}^{d})}.$
We can choose suitable $q_{1}$ and $q_{2}$ such that $p$ lies between $q_{1}$ and $q_{2}$. Interpolating these two bilinear estimates (26) and (27), there exists $s<p$ such that
@@
$\left\|\mathcal{E}_{\alpha}f_{\mathcal{A}_{0}^{j}}\right\|_{L^{q}(\mathbb{R}^{d+1})}^{2}\leq\left\|\sum_{\tau\sim\tau^{\prime}\subset\mathcal{A}_{0}^{j}}\mathcal{E}_{\alpha}f_{\tau}\mathcal{E}_{\alpha}f_{\tau^{\prime}}\right\|_{L^{\frac{q}{2}}(\mathbb{R}^{d+1})}\lesssim\left(\sum_{\tau\sim\tau^{\prime}\subset\mathcal{A}_{0}^{j_{0}}}\left\|\mathcal{E}f_{\tau}\mathcal{E}f_{\tau^{\prime}}\right\|_{L^{\frac{q}{2}}(\mathbb{R}^{d+1})}^{\theta_{1}}\right)^{\frac{1}{\theta_{1}}},$
where $\theta_{1}:=(q/2)_{*}>p/2>s/2$ and recall $(q/2)_{*}:=\min\{q/2,(q/2)^{\prime}\}$. By (28), we further have
@@
$\left\|\mathcal{E}_{\alpha}f_{\mathcal{A}_{0}^{j}}\right\|_{L^{q}(\mathbb{R}^{d+1})}^{2}\lesssim\left(\sum_{k}\sum_{\tau\in\mathcal{D}_{k}}|\tau|^{-2\theta_{1}(\frac{1}{s}-\frac{1}{p})}\|f_{\tau}\|_{L^{s}(\mathbb{R}^{d})}^{2\theta_{1}}\right)^{\frac{1}{\theta_{1}}}.$
@@
$\lesssim$ $\left(\sum_{k}\sum_{\tau\in\mathcal{D}_{k}}|\tau|^{-2\theta_{1}(\frac{1}{s}-\frac{1}{p})}\|f_{\tau}\|_{L^{s}(\mathbb{R}^{d})}^{2\theta_{1}}\right)^{\frac{1}{2\theta_{1}}}$
$\lesssim$ $\sup_{k\in\mathbb{Z}}\sup_{\tau\in\mathcal{D}_{k}}\sup_{l\geq 0}2^{-c_{1}l}\|f_{\tau,l}\|_{L^{p}(\mathbb{R}^{d})}^{\theta_{2}}\|f\|_{L^{p}(\mathbb{R}^{d})}^{1-\theta_{2}}$
for some $c_{1}>0$ and $10<\theta_{2}<$ 1. Then by rescaling, we obtain
$\left\|\mathcal{E}_{\alpha}f_{\mathcal{A}_{N}}\right\|_{L^{q}(\mathbb{R}^{d+1})}\lesssim\sup_{k\in\mathbb{Z}}\sup_{\tau\in\mathcal{D}_{k}}\sup_{l\geq 0}2^{-c_{1}l}\|f_{\tau,l}\|_{L^{p}(\mathbb{R}^{d})}^{\theta_{2}}\|f\|_{L^{p}(\mathbb{R}^{d})}^{1-\theta_{2}}$
holds for all $f\in L^{p}(\mathbb{R}^{d})$ with $\mathrm{supp}f\subset\mathcal{A}_{N}$. Combining this estimate and Lemma 4.1, we get the desired estimate (25) and finish the proof. ∎
@@
We are going to show that the desired decreasing sequence can be taken as $\rho_{j}:=C_{*}(j/2)^{-\theta/p}$, where $\theta$ is taken from Lemma 4.2 and $C_{*}$ is the optimal constant of the inequality (25). Then for fixed $j$, we investigate the following non-zero item
$\sup\limits_{\tau\in\mathcal{D}}\|f\chi_{\tau}\chi_{\{\xi:|f(\xi)|<2^{j}|\tau|^{-{1}/{p}}\|f\|_{L^{p}(\mathbb{R}^{d})}\}}\|_{L^{p}(\mathbb{R}^{d})}.$
Let the cube sequence $\tau_{n}=c_{n}+h_{n}[0,1]^{d}$ approach this supremum (bigger than half of this supremum). Then one can use the dominated convergence theorem to show that the cases $h_{n}\to\infty$ and $h_{n}\to 0$ cannot occur, and moreover, the case $c_{n}\to\infty$ cannot occur. These imply the fact that the aforementioned supremum can be reached.
@@
One can observe that the functions $f^{j}$ in Lemma 4.3, even with small support, may have large values, and conversely, $f^{j}$ with small values may have wide support. Next, we aim to control the support and the value simultaneously.
@@
If this proposition were to fail, for any dilation parameters $\left\{h_{n}\right\}\subset\mathbb{R}_{+}$, there would exist $\varepsilon>0$ and $\lim_{l\to\infty}m_{l}=\infty$ such that
@@
Noticing that there exists $J_{*}$ such that $\rho_{J_{*}}<M_{\alpha,p}$, thus up to subsequences, there exists one index $j_{0}\in\{1,2,\ldots,J_{*}\}$ such that
(33) $\|\mathcal{E}_{\alpha}f_{n}^{j_{0}}\|_{L^{q}(\mathbb{R}^{d+1})}\geq\frac{M_{\alpha,p}-\rho_{J_{*}}}{2J_{*}}>0,\quad\forall n\in\mathbb{N}.$
Without loss of generality, we may assume $j_{0}=1$ and then introduce the following notations
$\tau_{n}^{1}:=\xi_{n}^{1}+2^{-k_{n}^{1}}[0,1]^{d},\quad g_{n}^{1}(\xi):=2^{\frac{k_{n}^{1}d}{p}}f_{n}^{1}\left(2^{-k_{n}^{1}}\xi\right),\quad\Rightarrow\quad\left|g_{n}^{1}\left(\xi-2^{k_{n}^{1}}\xi_{n}^{1}\right)\right|\leqslant 2\chi_{[0,1]^{d}}(\xi).$
If there holds $2^{k_{n}^{1}}|\xi_{n}^{1}|\to\infty$ as $n\to\infty$, recalling the stationary phase consequence (6), then the fact $\alpha>2$ and dominated convergence theorem imply that
@@
$=\left(2^{k_{n}^{1}}|\xi_{n}^{1}|\right)^{2-\alpha}\int_{\mathbb{R}^{d+1}}\left|\int_{\mathbb{R}^{d}}e^{ix\xi+it\frac{|\xi-2^{k_{n}^{1}}\xi_{n}^{1}|^{\alpha}-\left(2^{k_{n}^{1}}|\xi_{n}^{1}|\right)^{\alpha}+\alpha\left(2^{k_{n}^{1}}|\xi_{n}^{1}|\right)^{\alpha-2}2^{k_{n}^{1}}\xi\cdot\xi_{n}^{1}}{\left(2^{k_{n}^{1}}|\xi_{n}^{1}|\right)^{\alpha-2}}}g_{n}^{1}\left(\xi-2^{k_{n}^{1}}\xi_{n}^{1}\right)\mathrm{d}\xi\right|^{q}\mathrm{d}x\mathrm{d}t$
$=o_{n\to\infty}(1),$
which contradicts the inequality (33). Hence, up to subsequences, there exists $|\xi^{1}|<\infty$ such that $2^{k_{n}^{1}}\xi_{n}^{1}\to\xi^{1}$ as $n\to\infty$. Thus by applying scaling symmetries, we can assume $\tau_{n}^{1}=\xi^{1}+[0,1]^{d}$ for all $n$ and the remainder cubes can be written as
@@
Then by our hypothesis, for such scaling symmetries $\{h_{n}\}$ we can get $\varepsilon$ and $m_{l}$ satisfying the inequality (31); and by taking $J$ large enough, the inequality (32) can guarantee that
@@
Next we are going to apply the generalized Brézis-Lieb lemma [20, Lemma 3.1], so we need to show that $\mathcal{E}_{\alpha}B_{n}^{J}(x,t)\to 0$ almost everywhere in $\mathbb{R}^{d+1}$ as $n\to\infty$. Since $J$ is a fixed finite number, we only need to investigate each item $\mathcal{E}_{\alpha}f_{n}^{j}$ with bad index $j$. For each fixed bad index $j$, we can divide the proof into three cases: if $k_{n}^{j}\to+\infty$ as $n$ goes to infinity then $f_{n}^{j}\to 0$ in $L^{1}(\mathbb{R}^{d})$, which by direct computation implies $\mathcal{E}_{\alpha}f_{n}^{j}\to 0$ almost everywhere; if $k_{n}^{j}\to-\infty$ as $n$ goes to infinity then $f_{n}^{j}\to 0$ in $L^{2}(\mathbb{R}^{d})$, which implies $\mathcal{E}_{\alpha}f_{n}^{j}\to 0$ in $L^{2+\frac{2\alpha}{d}}(\mathbb{R}^{d+1})$ and up to subsequences $\mathcal{E}_{\alpha}f_{n}^{j}\to 0$ almost everywhere; if $|k_{n}^{j}|$ remains bounded and $|\xi_{n}^{j}|\to\infty$ as $n$ goes to infinity, which implies $f_{n}^{j}\rightharpoonup 0$ in $L^{2}(\mathbb{R}^{d})$ and further the Lemma 2.4 implies $\mathcal{E}_{\alpha}f_{n}^{j}(t,x)\to 0$ almost everywhere. In summary, for fixed $J$, we obtain the desired pointwise convergence conclusion $\mathcal{E}_{\alpha}B_{n}^{J}\to 0$ almost everywhere. By the construction, for fixed $J$, there exists $G^{J}\in L^{p}(\mathbb{R}^{d})$ such that $G_{n}^{J}\to G^{J}$ in $L^{p}(\mathbb{R}^{d})$ as $n$ goes to infinity. Hence, by the aforementioned generalized Brézis-Lieb lemma and the fact $q>p$, we conclude